OpenStudy (anonymous):

Write an equation in point-slope form of the line that passes through the points (-3,-4) and (3,4).

OpenStudy (anonymous):

did you find the slope? or is it not clear how to do that?

OpenStudy (anonymous):

if it is not clear, say so, and i can show you it it not too hard

OpenStudy (anonymous):

\[m=\frac{ 4-(-4) }{ 3-(-3) }=\frac{ 8 }{ 6}\]

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Nnesha @Princer_Jones @jojo4eva

OpenStudy (anonymous):

ok good you are almost done

OpenStudy (anonymous):

btw \[\frac{8}{6}=\frac{4}{3}\] use the "point slope" formula and get \[y-4=\frac{4}{3}(x-3)\]

OpenStudy (anonymous):

solve for \(y\) and get \[y-y=\frac{4}{3}x-4\\ y=\frac{4}{3}x\]

OpenStudy (anonymous):